I have 2 numbers.The product of the two numbers equals 175.The sum of the two numbers is equal to 40.What are the two numbers?
Answer by checkley79(3341)

XY=175
X+Y=40
X=40-Y REPLACE X IN THE FIRST EQUATION BY(40-Y).
(40-Y)Y=175
40Y-Y^2=175
Y^2-40Y+175=0
(y-35)(y-5)=0
y-35=0
y=35 ans.
y-5=0
y=5 ans.
x*35=175
x=175/35=5 ans. when y=35.
x*5=175
x=175/5=35 ans. when y=5.